OpenStudy (anonymous):

will give medal Solve for x: 2|2x – 2| + 4 = 20 A. x = 5, x = –3 B. x = 5, x = –5 C. x = 3, x = –3 D. x = –5, x = 3

OpenStudy (anonymous):

subtract 4 from both sides: 2|2x-2| = 16 divide by two on both sides: |2x-2| = 8 now use casework and split the absolute value into 2 cases: 2x-2 = 8 or 2x-2 = -8 solving for both equations: x = 5, x = -3 the answer is A

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

I have one more if you don't mind

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what is it

OpenStudy (anonymous):

Solve for x: 3x – 2 > 5x + 10 A. x > 6 B. x < 6 C. x < –6 D. x > –6

OpenStudy (anonymous):

add 2 to both sides: 3x > 5x + 12 subtract 5x from both sides: -2x > 12 divide by two and switch the "greater than" to "less than": x < -6 the answer is C

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

you're welcome